OpenStudy (anonymous):

Solve. x2 + x = 20 Choices are: x = 5 and x = –4 x = 5 and x = 4 x = –5 and x = –4 x = –5 and x = 4

OpenStudy (abdela25):

x=-5 and x=4 thats your answer

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (abdela25):

yeah no problem but do you know how i got that becuase its good to know how the answer is found

OpenStudy (anonymous):

Not a clue

OpenStudy (abdela25):

yeah okay ill help out and maybe @arabpride will help me out with some explaining too :)

OpenStudy (arabpride):

XD then how did u get ur answer @abdela25 XD

OpenStudy (abdela25):

wellllll ill give you step by step lol

OpenStudy (arabpride):

ill do that as well..

OpenStudy (arabpride):

i dont think Martin cares for an explanation anymore..lol

OpenStudy (abdela25):

1. first i took one half of the coefficient of x and square it then add it too both sides add 1/4 to both sides \[x^2+x+1/4=81/4 \] 2. factor the left hand side...write the left hand side as a square \[(x+1/2)^2=81/4 \] eliminate the exponent on the left hand side...take the square root of both sides \[x+1/2=9/2 \] or \[x+1/2=-9/2\] 3. look at the first equation: solve for x subtract 1/2 from each side answer would be x=4 or x=5 what do you think @arabpride xD lol

OpenStudy (arabpride):

OHH i see what u did there..Good job!! *^.^*

OpenStudy (abdela25):

lol thankss ! Arab brain thoo thats all lol

OpenStudy (arabpride):

IKR!!! XD WE're takin ovaaaa!!! Masha-Allah!! We're SMART!!! XD lol

OpenStudy (arabpride):

*high five* NO actually im not...... Imma PROUD MUSLIM!!!!

OpenStudy (arabpride):

XD neither.. but Insh-Allah Peacce in both!!! IM from Yemen, AND DAMN Proud too!!

OpenStudy (arabpride):

possibly****

OpenStudy (abdela25):

yeahh good idea im gonna delte mine and you delete yours and you can msg me anytime thoo @arabpride

OpenStudy (arabpride):

alright bro!! i mean im not afraid of ppl seein ma postsor anythin, cuz if i can say it right up to pplz faces, i wouldnt mind puttin it on the internet..but its for the best..

OpenStudy (arabpride):

all done...lol

OpenStudy (abdela25):

yeah true true all that so yeahh okayy

OpenStudy (arabpride):

TTYL *^.^*

OpenStudy (abdela25):

okay see ya

OpenStudy (arabpride):

feel free to Pm me anytime, any day.. *^.^*